OpenStudy (anonymous):

: (a) ½mv2 = ½mv02 + √mgh, (b) v = v0 + at2, (c) ma = v2. . Do a dimensional analysis of each equation and explain why the equation can’t be correct.

OpenStudy (anonymous):

kg* (m/s)^2 = kg*(m/s)^2+sqrt[kg*m/s^2 *m] kg* m^2/s^2= kg *m^2/s^2+ sqrt[kg] m/s ^ ^ unit for these two don't match, thus you can't add them;

OpenStudy (anonymous):

Ok thank you.